HELP!!!! Here's my problem that escapes me:
When the digits of a two-digit number are reversed, the new number is 9 more than the original number, and the sum of the digits of the original number is 11. What is the original number?
Thanks in advance for your help.
Joanne

Answer by scott8148(6628)

let x="tens digit" and y="units digit"

"When the digits of a two-digit number are reversed, the new number is 9 more than the original number"
__ 10y+x=10x+y+9

"the sum of the digits of the original number is 11"
__ x+y=11

x=11-y __ substituting gives 10y+(11-y)=10(11-y)+y+9 __ 9y+11=119-9y __ 18y=108